# Estimating the FDR of variable selection

### v2

@@ 3 Standard error of $\widehat{\textnormal{FDR}}$
In this section, we discuss the standard error of our estimator $\widehat{\textnormal{FDR}}$. Since our general method can be adapted for many different selection algorithms in many different statistical modeling contexts, it is very difficult to provide a general analysis that covers all contexts in which our method might be applied, but we can offer theoretical results in a narrower context, as well as a practical estimation strategy that we have found to be reasonably reliable in our simulation studies. Section 3.1 develops a theoretical finite-sample bound for our estimator under stylized assumptions, leading to conditions under which the variance of $\widehat{\textnormal{FDR}}$ is vanishing, while Section 3.2 discusses a specialized bootstrap method for standard error estimation, which we use to produce all figures in this work.
@@ 3.1 Theoretical bound for standard error
Let $\boldsymbol{X}$ be an $n\times d$ matrix, and let $B_{1},\ldots,B_{K}\subseteq[d]$ represent a partition of $[d]$ into $K$ disjoint subsets (that is, $\bigcup_{k}B_{k}=[d]$). We say the matrix $\boldsymbol{X}$ has block orthogonal variables with respect to the partition $B$ if $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{\ell}}=\boldsymbol{0}_{|B_{k}|\times|B_{\ell}|}$, for all $k\neq\ell$.
Assumption 3.1 allows us to treat the Lasso on the entire matrix $\boldsymbol{X}$ as a separable optimization problem on the $K$ different blocks of explanatory variables. When this assumption is in effect, we will denote the size of the largest block as $b=\max_{k}|B_{k}|$. We will assume throughout this section that every explanatory variable has mean zero and unit norm: $\sum_{i=1}^{n}x_{i,j}=0$ and $\|\boldsymbol{X}_{j}\|^{2}=1$, for all $j=1,\ldots,d$. This last assumption comes with little practical loss of generality since conventional implementations of the Lasso begin by normalizing all variables in this way.
For our results below, it will be important to be able to provide a probabilistic lower bound on the average number $q$ of selections per block:
@@
We have $\mathbb{P}(R\leq Kq)\leq\rho$ for some $q,\rho\geq 0$.
@@
    Let $\delta(z):=2(1-\Phi(z))$ denote the probability that a Gaussian random variable with mean 00 and variance 1 exceeds $z$ in absolute value. Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\delta(\lambda/\sigma))$ random variable, and for any $q<\delta(\lambda/\sigma)$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\delta(\lambda/\sigma)-q)^{2}}{2\delta}\cdot K\right).$
@@
    Assume no fewer than $K\xi$ blocks contain at least one variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$ for some $\xi\in(0,1]$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$. Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\xi/2)$ random variable, and for any $q<\xi/2$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
A finite-sample upper bound for $\text{Var}(\widehat{\textnormal{FDR}})$ for Lasso with block-orthogonal variables then follows.
@@
$\textnormal{Var}(\widehat{\textnormal{FDR}})\leq O\left(\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}+\frac{b^{2}}{q^{2}}\cdot\frac{1}{n-d}+b^{4}K^{2}(\rho\vee e^{-(n-d)/32})\right),$
@@
    Conditioning on $\boldsymbol{S}_{j}$ fixes $\boldsymbol{X}_{k}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ for every $k$ that is not in the same block as $j$. As a result, resampling the data conditional on $\boldsymbol{S}_{j}$ would not change the selection in all other blocks. Hence, the number of selections is conditionally stable: with resampling conditional on $\boldsymbol{S}_{j}$ it is $R(\boldsymbol{D})+O(b)$.
@@
    For any $j\in B_{k}$, let $X_{B_{k}-j}$ denotes the variables in block $B_{k}$ except $X_{j}$. The selection of variable $X_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Formally, we have
    $E_{j}:=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]\approx\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{S}_{j}].$
    Note these $E_{j}$ are block-independent. [^5] Alert readers might have noticed that conditioning on $\boldsymbol{S}_{j}$ is not the same as conditioning on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$: there is an additional $\left\|\boldsymbol{Y}\right\|$ term. The analysis of the effects of $\left\|\boldsymbol{Y}\right\|$ is rather technical and it is the source of the $O(1/(n-d))$ term in the variance bound.
@@
    Likewise, the p-value $p_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. So $\phi_{j}(p_{j})$ are block-independent.
@@
$\widehat{\textnormal{FDR}}_{j}\;=\;\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\phi_{j}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R+O(b)}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R}.$
The last approximation holds when $R>Kq$ with high probability and $K$ is substantially larger than $b$. Then
$\widehat{\textnormal{FDR}}\;=\;\sum_{j}\widehat{\textnormal{FDR}}_{j}\;\approx\;\frac{\sum_{k}\left(\sum_{j\in B_{k}}E_{j}\cdot\phi_{j}\right)}{R}.$
This is an average of $K$ independent random variables. Therefore, the variance is of order $O(1/K)$ by the concentration laws. A formal asymptotic result is shown in Corollary 3.1.
@@
Under the same settings as in Theorem 3.1, consider a series of problems where we hold $c_{\phi}$, $b$, $L_{\psi}$ and $\sigma$ fixed but allow $\boldsymbol{\theta}$, $\lambda$, and $n$ to vary as $K$ increases. Suppose that at least $s(K)$ blocks contain a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
@@
    For $40<\varepsilon<1/$ 4, suppose $s(K)\geq K^{1-4\varepsilon}$, and let $K\to\infty$ with $n-d\geq K^{1-2\varepsilon}$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1+4\varepsilon})\to 0.$
@@
    For $10<\xi<$ 1, suppose $s(K)\geq K\xi$, and let $K\to\infty$ with $n-d\geq 64\log(K)$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1})\to 0.$
@@ 3.2 Standard error estimation by bootstrap
To ameliorate this issue, we try to construct $\widehat{P}$ in such a way that most of the null variables in $P$ remain null in $\widehat{P}$, but the signal variables are estimated with little bias. Formally, let $\widehat{\mathcal{H}_{0}}$ denote an estimate of the index set of noise variables. In the parametric setting $\boldsymbol{D}\sim P_{\boldsymbol{\theta}}$, we find a sparse MLE of $\boldsymbol{\theta}$, denoted as $\hat{\boldsymbol{\theta}}$, under the constraints that $\theta_{j}=0$ for all $j\in\widehat{\mathcal{H}_{0}}$. Then we set $\widehat{P}=P_{\hat{\boldsymbol{\theta}}}$ and draw bootstrap samples to calculate $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$. We formally describe our bootstrap scheme in the parametric setting in Algorithm 1.
@@
$\hat{\boldsymbol{\theta}}(\widehat{\mathcal{H}_{0}};\boldsymbol{D})=\underset{\theta_{j}=0,\,j\in\widehat{\mathcal{H}_{0}}}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D})$
@@
  generate sample $\hat{\boldsymbol{D}}^{(m)}\sim P_{\hat{\boldsymbol{\theta}}}$;  compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
@@
In the Gaussian linear model, the negative log-likelihood $-\log P_{\theta}(\boldsymbol{D})$ is proportional to the mean-squared error. Therefore, $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}})$ is the least square coefficients regressing $Y$ on variables $X_{j}$ for $j\in\widehat{\mathcal{H}_{0}}^{\mathsf{c}}$.
@@
For the model-X setting, we can employ an analogous method by using a variant of the i.i.d. bootstrap that constrains $H_{j}:\;Y\perp\!\!\!\perp X_{j}\mid X_{-j}$ to be true for all $j\in\widehat{\mathcal{H}_{0}}$. Specifically, we first draw bootstrap samples of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$ from the empirical distribution of the data. Then, we sample $X_{\widehat{\mathcal{H}_{0}}}$ from the conditional distribution $X_{\widehat{\mathcal{H}_{0}}}\mid(Y,X_{-\widehat{\mathcal{H}_{0}}})$, which is equal to $X_{\widehat{\mathcal{H}_{0}}}\mid X_{-\widehat{\mathcal{H}_{0}}}$ under the intersection null $\bigcap_{j\in\widehat{\mathcal{H}_{0}}}H_{j}$. We formally describe our bootstrap scheme in the model-X nonparametric setting in Algorithm 2.
@@
Preparation: For simplicity, dentoe $x_{i,\widehat{\mathcal{H}_{0}}}=(x_{i,j},\;j\in\widehat{\mathcal{H}_{0}})$ as the $i$th observation of $X_{\widehat{\mathcal{H}_{0}}}$. Let $\widehat{P}_{-}$ denote the empirical distribution of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$
$\widehat{P}_{-}(\,\cdot\,)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{(y_{i},x_{i,-\widehat{\mathcal{H}_{0}}})\in\,\cdot\,\right\}.$

### v4

@@ 3 Standard error of $\widehat{\textnormal{FDR}}$
In this section, we discuss the standard error of our estimator $\widehat{\textnormal{FDR}}$. Since our general method can be adapted for many different selection algorithms in many different statistical modeling contexts, it is very difficult to provide a general analysis that covers all contexts in which our method might be applied, but we can offer theoretical results in a narrower context, as well as a practical estimation strategy that we have found to be reasonably reliable in our simulation studies. Section 3.1 develops a theoretical finite-sample bound for our estimator under stylized assumptions, leading to conditions under which the variance of $\widehat{\textnormal{FDR}}$ is vanishing. Section 3.2 discusses a specialized bootstrap method for standard error estimation, which we use to produce all figures in this work. Section 3.3 introduces the theoretical property of our bootstrap scheme.
@@ 3.1 Theoretical bound for standard error
Let $\boldsymbol{X}$ be an $n\times d$ matrix, and let index set $(b)\subseteq[d]$ for $b=1,\ldots,B$ represent a partition of $[d]$ into $B$ disjoint subsets (that is, $\bigcup_{b}(b)=[d]$). For each $b$, let $\boldsymbol{X}_{(b)}\in\mathbb{R}^{n\times m_{b}}$ be the submatrix of $\boldsymbol{X}$ containing only the columns corresponding to the variables in $(b)$, where $m_{b}=|(b)|$ is size of block $(b)$. We say that $\boldsymbol{X}$ has block orthogonal variables with respect to the partition $\{(b)\}_{b=1}^{B}$ if
$\boldsymbol{X}_{(b_{1})}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b_{2})}=\boldsymbol{0}_{m_{b_{1}}\times m_{b_{2}}},\quad\text{for all }b_{1}\neq b_{2}.$
Assumption 3.1 allows us to treat the Lasso on the entire matrix $\boldsymbol{X}$ as a separable optimization problem on the $B$ different blocks of explanatory variables. When this assumption is in effect, we will denote the size of the largest block as $m=\max_{b}m_{b}$. We will assume throughout this section that every explanatory variable has mean zero and unit norm: $\sum_{i=1}^{n}x_{i,j}=0$ and $\|\boldsymbol{X}_{j}\|^{2}=1$, for all $j=1,\ldots,d$. This last assumption comes with little practical loss of generality since conventional implementations of the Lasso begin by normalizing all variables in this way.
For our results below, it will be important to be able to provide a probabilistic lower bound $q$ on the average number of selections per block:
@@
We have $\mathbb{P}(R\leq Bq)\leq\rho$ for some $q,\rho\geq 0$.
@@
    Let $\delta(z):=2(1-\Phi(z))$ denote the probability that a Gaussian random variable with mean 0 and variance 1 exceeds $z$ in absolute value. Then $R$ is stochastically larger than a $\textnormal{Binom}(B,\delta(\lambda/\sigma))$ random variable, and for any $q<\delta(\lambda/\sigma)$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\delta(\lambda/\sigma)-q)^{2}}{2\delta}\cdot B\right).$
@@
    Assume no fewer than $B\xi$ blocks contain at least one variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$ for some $\xi\in(0,1]$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$. Then $R$ is stochastically larger than a $\textnormal{Binom}(B\xi,1/2)$ random variable, and for any $q<\xi/2$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\xi-2q)^{2}}{4\xi}\cdot B\right).$
A finite-sample upper bound $\text{Var}(\widehat{\textnormal{FDR}}) for$ Lasso with block-orthogonal variables then follows.
@@
$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(\left(\frac{m^{4}}{q^{4}}+\frac{m^{6}}{Bq^{4}}\right)\cdot\frac{1}{B}+\frac{m^{2}}{q^{2}}\cdot\frac{1}{n-d}+m^{6}B^{2}(\rho\vee e^{-(n-d)/32})\right),$
@@
    Conditioning on $\boldsymbol{S}_{j}$ fixes $\boldsymbol{X}_{k}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ for every $k$ that is not in the same block as $j$. As a result, resampling the data conditional on $\boldsymbol{S}_{j}$ would not change the selection in all other blocks. Hence, the number of selections is conditionally stable: with resampling conditional on $\boldsymbol{S}_{j}$ it is $R(\boldsymbol{D})+O(m)$.
@@
    For any $j\in(b)$, let $\boldsymbol{X}_{(b);-j}$ denote the variables in block $(b)$ except $\boldsymbol{X}_{j}$. The selection of variable $\boldsymbol{X}_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Formally, we have
    $E_{j}:=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]\approx\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{S}_{j}].$
    Note these $E_{j}$ are block-independent. [^7] Alert readers might have noticed that conditioning on $\boldsymbol{S}_{j}$ is not the same as conditioning on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$: there is an additional $\left\|\boldsymbol{Y}\right\|$ term. The analysis of the effects of $\left\|\boldsymbol{Y}\right\|$ is rather technical and it is the source of the O(1/(nd))O(1/(n-d)) term in the variance bound.
@@
    Likewise, the p-value $p_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. So $\phi_{j}(p_{j})$ are block-independent.
@@
$\widehat{\textnormal{FDR}}_{j}\;=\;\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\phi_{j}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R+O(m)}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R}.$
The last approximation holds when $R>Bq$ with high probability and $B$ is substantially larger than $m$. Then
$\widehat{\textnormal{FDR}}\;=\;\sum_{j}\widehat{\textnormal{FDR}}_{j}\;\approx\;\frac{\sum_{k}\left(\sum_{j\in(b)}E_{j}\cdot\phi_{j}\right)}{R}.$
This is an average of $B$ independent random variables. Therefore, the variance is of order $O(1/B)$ by the concentration laws. A formal asymptotic result is shown in Corollary 3.1.
@@
Under the same settings as in Theorem 3.1, consider a series of problems where we hold $c_{\phi}$, $m$, $L_{\phi}$ and $\sigma$ fixed but allow $\boldsymbol{\theta}$, $\lambda$, and $n$ to vary as $B$ increases. Suppose that at least $s(B)$ blocks contain a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
@@
    For $0<\varepsilon<1/4$, suppose $s(B)\geq B^{1-4\varepsilon}$, and let $B\to\infty$ with $n-d\geq B^{1-2\varepsilon}$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(B^{-1+4\varepsilon})\to 0.$
@@
    For $0<\xi<1$, suppose $s(B)\geq B\xi$, and let $B\to\infty$ with $n-d\geq 64\log(B)$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(B^{-1})\to 0.$
@@ 3.2 Standard error estimation by bootstrap
[…]
@@
$\hat{\boldsymbol{\theta}}^{(*)}(\widehat{\mathcal{H}_{0}};\boldsymbol{D})=\underset{\theta_{j}=0,\,j\in\widehat{\mathcal{H}_{0}}}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D})$
@@
  generate sample $\hat{\boldsymbol{D}}^{(m)}\sim P_{\hat{\boldsymbol{\theta}}^{(*)}}$;  compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
@@
[…]
@@
[…]
@@
[…]
3.3 Theory of parametric bootstrap variance estimation
It is difficult to study the theoretical property of the bootstrap variance estimator in full generality. Nevertheless, we make substantial progress for Gaussian linear models with Lasso selection. In particular, we revisit the stylized block-orthogonal Gaussian linear model used in our theoretical variance analysis (Section 3.1) with known error variance $\sigma^{2}$ for simplicity and consider the Lasso selection with a appropriately chosen penalty level in an asymptotic regime where the true and estimated FDR are bounded away from 0 and $1$ in the limit.
For variance estimation, we consider the parametric bootstrap in Algorithm 1 with $M=\infty$, which assumes away Monte Carlo error, and the estimated null set
$\widehat{\mathcal{H}_{0}}=\{j:\ \widehat{\theta}^{\textnormal{Lasso}}_{j}(\bar{\lambda})=0\},$
where $\bar{\lambda}>\lambda$ is a slightly inflated penalty level used only for estimating the bootstrap model. We then choose $\hat{\boldsymbol{\theta}}^{(*)}$ as the OLS estimator restricting to $\widehat{\mathcal{H}_{0}}$, as described in Example 3.1. We call the resulting model an OLS-after-Lasso model. Define $\widehat{\textnormal{FDR}}^{(*)}$ as a generic bootstrap sample of $\widehat{\textnormal{FDR}}$ computed under OLS-after-Lasso model.
In this setting, we prove that the asymptotic variance of the parametric bootstrap FDR estimates conditional on data matches the asymptotic variance of our FDR estimator. We elaborate the setting and technical details in Appendix E.1.
Theorem 3.2.
Consider a Gaussian linear model with known $\sigma^{2}$ that satisfies the Assumption 3.1. In the setting of Theorem E.2 in Appendix E.2, there exist deterministic sequences $\nu_{d}=o(1)$ and $\tilde{\alpha}_{d}$ that are bounded away from 0 and $1$, and a data-dependent sequence $\alpha_{d}^{(*)}=\alpha_{d}^{(*)}(\boldsymbol{X},\boldsymbol{Y})$ such that
$\sup_{t\in\mathbb{R}}\biggl|\mathbb{P}\!\biggl(\frac{1}{\nu_{d}}\bigl(\widehat{\textnormal{FDR}}-\tilde{\alpha}_{d}\bigr)\leq t\biggr)-\Phi(t)\biggr|=o(1),$
and
$\sup_{t\in\mathbb{R}}\biggl|\mathbb{P}\!\biggl(\frac{1}{\nu_{d}}\bigl(\widehat{\textnormal{FDR}}^{(*)}-\alpha_{d}^{(*)}\bigr)\leq t\,\Big|\,\boldsymbol{X},\boldsymbol{Y}\biggr)-\Phi(t)\biggr|=o_{\mathbb{P}}(1),$
where $\Phi$ denotes the CDF of $\mathcal{N}(0,1)$.